OpenStudy (anonymous):

solve the system of equations using substitution. x^2 + y^2 = 61 x + y = 11

OpenStudy (anonymous):

Solve for one of the variables in x+y=11

OpenStudy (anonymous):

Then plug that in for one of the variables in the other equation

OpenStudy (anonymous):

i have to use substitution. so i have to plug one of them into the other.

OpenStudy (anonymous):

...that's what I said :3 That's what substitution is. Say you want to find x first. Solve for y in x+y=11, which would be y=11-x, and substitute that for y, then solve for x.

OpenStudy (anonymous):

You would get x^2+(11-x)^2=61

OpenStudy (anonymous):

Solve that for x, then use x and solve for y in the same equation.

OpenStudy (tkhunny):

Well, do it. y = 11 - x \(x^{2} + (11-x)^{2} = 61\) That's about as simple as any other Quadratic Equation. Let's see what you get. Keep in mind a circle and a line can produce three possible results: 1) No intersection at all. 2) One tangent intersection. 3) Two intersections.

OpenStudy (anonymous):

ok wait :)

OpenStudy (anonymous):

yeah well i know its going to be two intersections :)

OpenStudy (tkhunny):

Fair enough, but they are a little funny and the casual observer may think there is only one.

OpenStudy (anonymous):

no, the solutions are (6,5) (5,6)

OpenStudy (tkhunny):

Good call. Their reflection on x = y can be confusing. I am glad to see you were not confused.

Directrix (directrix):

x^2 + y^2 = 61 x + y = 11 ------------ By substitution. x + y = 11 y = 11 - x

Directrix (directrix):

Substituting (11 - x) for y in --> x^2 + y^2 = 61

Directrix (directrix):

x^2 + (11 - x) ^2 = 61 x^2 + 121 -22x + x^2 = 61

OpenStudy (anonymous):

okaii got all that until now:)

Directrix (directrix):

2 x^2 - 22x + 121 = 61 2x^2 -22x + 60 = 0

Directrix (directrix):

x^2 - 11x + 30 = 0 (x - 6) ( x - 5) = 0 x - 6 = 0 OR x - 5 = 0 x = 6 OR x = 5

Directrix (directrix):

x = 6 OR x = 5 -------------- Let x = 6. Then, y = 11 - x y = 11 - 6 y = 5 One set of coordinates for one point of intersection is (6,5) -->

Directrix (directrix):

Let x = 5 --------- y = 11 - x y = 11 - 5 = 6. The second set of coordinates for the second point of intersection is (5, 6) -->

OpenStudy (anonymous):

thanx man :)))))

Directrix (directrix):

Glad to help.